OpenStudy (anonymous):

Which Pair Of Points Defines A Line With A Slope Of -4/3

OpenStudy (anonymous):

The Anwer Has To Be Out Of (-1, 6) and (-4, 10) B. (6, -1) and (-4, 10) C. (-1, 6) and (10, -4) D. (6, -1) and (10, -4)

OpenStudy (anonymous):

Help @Ganpat

OpenStudy (ganpat):

slope (m) = rise/run = (y2 - y1)/ (x2 - x1)... now can u check, with this formula ??

OpenStudy (anonymous):

iM Thinking D. What Do You Think Ganpat

OpenStudy (ganpat):

ha ha, u r smart dude.. LOL

OpenStudy (ganpat):

-4/3 = (y2 - y1)/ (x2 - x1) A = -4/3 B = - 11/10 C = - 10/ 11 D = -3/4.. Now, u try..

OpenStudy (anonymous):

Ganpat Try This 1 On 4 Size Lol Which of these pairs of points defines a line with a slope of -1? A. (3, 10) and (3, 5) B. (3, 10) and (5, 3) C. (10, 3) and (15, -2) D. (10, 3) and (3, 5)

OpenStudy (ganpat):

i tried.. now its ur turn.. :-p.. LOl..

OpenStudy (anonymous):

A. (3, 10) and (3, 5)=0,-5 B. (3, 10) and (5, 3)=-2,7 C. (10, 3) and (15, -2)=-5,5 D. (10, 3) and (3, 5)=7,-2 Am iRight? Lol

OpenStudy (ganpat):

well, the answer is C... (-2-3)/ (15-10 ) = -5/5 = -1...

OpenStudy (anonymous):

Yes iGot iT

OpenStudy (ganpat):

u r learning fast... Gooooooooddd !!!

OpenStudy (anonymous):

But What iF iT iN Reverse Like....... Define A Line A Slope Of 2/3 A. (2, 6) and (3, 0) B. (6, 2) and 3, 0) C. (6, 2) and (0, 0) D. (2, 6) and (0, 0)

OpenStudy (anonymous):

A. (2, 6) and (3, 0)=-1,6 B. (6, 2) and 3, 0)=3,2 C. (6, 2) and (0, 0)=6,2 D. (2, 6) and (0, 0)=2,6